# DreamTalk: When Expressive Talking Head Generation Meets Diffusion Probabilistic Models

### v1

@@ Abstract
Diffusion models have shown remarkable success in a variety of downstream generative tasks, yet remain under-explored in the important and challenging expressive talking head generation. In this work, we propose a DreamTalk framework to fulfill this gap, which employs meticulous design to unlock the potential of diffusion models in generating expressive talking heads. Specifically, DreamTalk consists of three crucial components: a denoising network, a style-aware lip expert, and a style predictor. The diffusion-based denoising network is able to consistently synthesize high-quality audio-driven face motions across diverse expressions. To enhance the expressiveness and accuracy of lip motions, we introduce a style-aware lip expert that can guide lip-sync while being mindful of the speaking styles. To eliminate the need for expression reference video or text, an extra diffusion-based style predictor is utilized to predict the target expression directly from the audio. By this means, DreamTalk can harness powerful diffusion models to generate expressive faces effectively and reduce the reliance on expensive style references. Experimental results demonstrate that DreamTalk is capable of generating photo-realistic talking faces with diverse speaking styles and achieving accurate lip motions, surpassing existing state-of-the-art counterparts.
1 Introduction
Audio-driven talking head generation, which concerns animating portraits with speech audio, has garnered significant interest due to its diverse applications in video games, film dubbing, and virtual avatars. Generating lifelike facial expressions is essential for enhancing the realism of talking heads [19]. These expressions during speech are termed as speaking styles [12, 46]. GANs [21] currently hold the state-of-the-art in expressive talking head generation [37, 30, 46, 20].However, their inherent issues with mode collapse and unstable training hamper their efficacy in consistently achieving high performance across a diverse range of speaking styles. Another issue is that prior methods often rely on reference videos [37, 30, 78] or texts [45, 90] to specify speaking styles. Their acquisition requires extra manual effort and hence is inflexible.
@@ 1 Introduction
Figure 1: Leveraging the powerful diffusion models, DreamTalk is able to generate highly expressive talking heads across diverse speaking styles. Furthermore, DreamTalk is able to derive personalized speaking style directly from input audio, which obviates the need for additional style references.
@@
As a new line of generative technique, diffusion models [26, 62] have recently been shown to produce high-quality results in numerous generative areas such as image generation [16, 53], video generation [60, 86, 27], and human motion synthesis [1, 72]. The success of diffusion models, stemming from their superior properties such as powerful distribution learning [16, 72], good convergence, and stylistic diversity, make them exceptionally promising for exploring expressive talking head generation. However, current diffusion-based talking head approaches [58, 66, 96] primarily concentrate on generating talking heads with neutral expressions and still struggle to produce satisfactory performance, e.g., suffering from frame jittering problem [58]. As a result, how to stimulate the full potential of diffusion models for expressive talking head generation is a promising yet untapped research direction.
@@
In this paper, we propose DreamTalk, an expressive talking head generation framework that takes advantage of diffusion models to simultaneously deliver high performance across diverse speaking styles and reduce the reliance on expensive style references. Specifically, DreamTalk is composed of a denoising network, a style-aware lip expert, and a style predictor. The diffusion-based denoising network produces audio-driven facial motions with the speaking style specified by a reference video. The great distribution-learning characteristic of diffusion models endows the denoising network with the potential to produce high-quality results across diverse speaking styles. To harness this potential, we design a style-aware lip expert that drives the denoising network to produce accurate lip motions with vivid expressions. Contrasting with previous lip experts that overlook expression information and thus compromise style expressiveness, the proposed lip expert not only enhances lip accuracy but also ensures expressiveness. Finally, to further eliminate the need for additional style references, a diffusion-based style predictor is incorporated to predict personalized speaking styles directly from audio. The style predictor also incorporates the portrait as input during prediction, leveraging the correlation between speaker identity and speaking styles, thereby enhancing performance.
Ultimately, DreamTalk can consistently generate photo-realistic talking faces with precise lip-sync across a wide range of speaking styles while minimizing the need for additional style references. It also enables versatile manipulation of speaking styles and exhibits robust generalization across varied inputs, including songs, speech in multiple languages, noisy audio, and out-of-domain portraits. The effectiveness of DreamTalk is demonstrated through comprehensive qualitative and quantitative evaluations, showcasing its superiority over existing state-of-the-art methods.
2 Related Work
@@
Audio-Driven Talking Head Generation. Audio-driven methods [14, 22, 81, 102, 67, 80] fall into two main categories: person-specific and person-agnostic. Person-specific approaches [68, 18, 29, 82, 41] are constrained to generating videos for speakers seen during training. Many of these [94, 74, 64, 35, 34, 29, 97, 98, 70] first craft 3D facial animations, later converting them into realistic videos. Recent advancements [23, 39, 57, 71, 93, 71] have employed neural radiance fields for modeling, yielding high-fidelity, realistic videos. Conversely, person-agnostic methods [10, 84, 55, 77] target generating videos for unseen speakers. Early methods prioritized lip synchronization [65, 5, 103, 6, 77, 49]. Later works shifted focus to natural facial expressions [95, 105] and head poses [7, 101, 83, 104, 100].
@@
Expressive Talking Head Generation. Early methods [90, 20, 13, 24, 69, 82, 29, 61] model expressions in discrete emotion classes. To model more fine-grained expressions, most recent methods [30, 37, 46] leverage an expression reference video and transfer the expressions from that video to the generated one. However, these GAN-based models suffer from mode collapse, leading to videos with inferior lip-sync and style expressiveness. Our work addresses these issues by using diffusion models.
@@
Specifying desired speaking styles effortlessly is also important for users. Most previous methods specify speaking styles using reference videos [30, 37, 46] or text [45, 90, 20], which needs human labor. A more user-friendly approach is to derive speaking styles from the input audio. Previous methods can only infer a limited number of discrete emotion classes from audio signals [29, 90, 61]. TH-PAD [95] generates expressions only aligned with the audio rhythm, not aligning with the emotional content of the audio. Besides, previous methods neglect the information in the input portrait. In this work, we aim to infer personalized and emotional expressions using input audio and portraits.
@@
Diffusion Models. Diffusion models [62, 26] have demonstrated strong performance across multiple vision tasks [16, 53, 86, 99], including text-to-image generation [54, 31], image inpainting [42, 33], human motion generation [1, 72], 3D content generation [43], and video generation [60, 50, 89, 87, 79]. Previous efforts [58, 66, 3, 47] employing diffusion models in talking head generation only produce talking heads with neutral emotion and the results are unsatisfactory. Some methods devise diffusion-based renderers [91, 17] or face motion prior [96], yet still use GAN or regression models to model the audio-motion mapping. In this work, we endeavor to harness diffusion models for the generation of expressive talking heads, presenting a more intricate challenge with greater practical relevance.
3 Method
@@
Figure 2: Illustration of DreamTalk. A style-aware lip expert (b), which evaluates the lip-audio synchronous probability under a given speaking style, is first trained to provide lip motion guidance for the denoising network (a). The denoising network is then trained to take the audio, the style reference video, and noisy face motion as input and predict the unnoised face motion. Then, A style predictor (c) is trained to predict the style code extracted from a video, taking audio and the speaker in that video as input. During inference process (d), the speaking style can be specified using style codes that are extracted from videos or derived from audio.
3.1 Problem Formulation
@@
Given a portrait $\bm{I}$, a speech $\bm{A}$, and a style reference video $\bm{R}$, our method aims to generate a talking head video with lip motions synchronized with the speech and the speaking style reflected in the reference video. The audio $\bm{A}=[\bm{a}_{i}]_{i=1}^{L}$ is parameterized as a sequence of acoustic features. $\bm{R}$ is a sequence of video frames.
@@
Besides, to eliminate the need for extra style references, our method also aims to infer the speaking style using solely the speech and the portrait. The inferred speaking style can replace the role of style reference videos in controlling the expressions, which enables our method to generate expressive talking head videos with only speech and portrait input.
3.2 DreamTalk
@@
DreamTalk comprises 3 key components: a denoising network, a style-aware lip expert, and a style predictor.
@@
The denoising network computes face motion conditioned on the speech and style reference video. The face motion $\bm{M}=[\bm{m}_{l}]_{l=1}^{L}$ is parameterized as a sequence of expression parameters from 3D Morphable Models [4]. The face motion is rendered into video frames by a renderer [52]. The style-aware lip expert provides lip motion guidance under diverse expressions and thus drives the denoising network to achieve accurate lip-sync while ensuring style expressiveness. The style predictor can predict the speaking style aligned with that conveyed in speech.
@@
The asterisk(***) indicates quantities that are generated.
@@
The style-aware lip expert encodes the lip motions and audio into respective embeddings conditioned on style reference and then computes the cosine similarity to represent the sync probability. To obtain lip motion information from face motion $\bm{m}$, we first convert $\bm{m}$ into the corresponding face mesh and select vertices in the mouth area as the representation of the lip motion [46]. The lip motion and audio encoders are mainly implemented by MLPs and 1D-convolutions, respectively. The style condition is fused into embeddings by first extracting style features from style reference using a style encoder, which mirrors the architecture of the one in the denoising network, and then concatenating the style features with intermediate feature maps from embedding encoders. The style encoder in the lip expert and the generator do not share parameters.
Style Predictor. Specifically, the style predictor $S_{\phi}$ predicts the style code $\bm{s}$ extracted by the style encoder in the trained denoising network. Observing the correlation between speaker identity and style codes (Sec. 4.4), the style predictor also integrates the portrait as input. The style predictor is instantiated as a diffusion model and is trained to predict the style code itself:
@@
MakeItTalk [105] 0.73 / 0.57 / 0.52 0.11 / 0.24 / 0.24 3.97 / 5.12 / 6.29 5.32 / 4.55 / 5.15 2.10 / 3.16 / 2.17
Wav2Lip [49] 0.80 / 0.63 / 0.54 0.18 / 0.30 / 0.30 2.72 / 4.53 / 5.85 4.05 / 3.60 / 4.64 5.26 / 5.83 / 5.70
PC-AVS [104] 0.50 / 0.42 / 0.36 0.07 / 0.13 / 0.09 5.83 / 9.71 / 12.9 4.97 / 4.17 / 7.42 2.18 / 4.85 / 4.73
AVCT [84] 0.83 / 0.76 / 0.64 0.14 / 0.22 / 0.23 2.92 / 2.86 / 3.62 5.52 / 3.57 / 3.71 2.53 / 4.27 / 3.89
GC-AVT [37] 0.34 / 0.36 /    - 0.14 / 0.28 /    - 8.04 / 10.2 /    - 7.10 / 6.23 /    - 2.42 / 4.72 /    -
EAMM [30] 0.40 / 0.40 / 0.43 0.08 / 0.14 / 0.20 6.70 / 7.03 / 6.36 6.48 / 6.86 / 4.89 1.41 / 2.54 / 2.24
StyleTalk [46] 0.84 / 0.81 / 0.66 0.16 / 0.30 / 0.29 2.12 / 1.96 / 2.92 3.25 / 2.41 / 2.96 3.47 / 4.82 / 4.51
SadTalker [100] 0.69 / 0.77 / 0.44 0.16 / 0.24 / 0.19 4.12 / 5.99 / 9.12 4.37 / 4.07 / 6.11 2.76 / 4.35 / 4.38
PD-FGC [78] 0.49 / 0.41 / 0.35 0.05 / 0.13 / 0.12 5.50 / 9.50 / 12.5 4.10 / 4.23 / 8.19 2.27 / 4.68 / 4.64
EAT [20] 0.53 / 0.59 / 0.47 0.15 / 0.26 / 0.20 5.54 / 3.86 / 5.53 4.79 / 4.03 / 5.88 2.16 / 4.54 / 4.35
DreamTalk 0.86 / 0.85 / 0.69 0.16 / 0.31 / 0.30 1.93 / 1.80 / 2.69 2.91 / 2.15 / 2.72 3.78 / 5.17 / 4.90
Ground Truth 1 / 1 / 1 0.22 / 0.31 / 0.33 0 / 0 / 0 0 / 0 / 0 4.13 / 5.44 / 5.23
Table 1: Quantitative comparisons on MEAD, HDTF, and Voxceleb2. Since we only receive GC-AVT samples on MEAD and HDTF, GC-AVT is not evaluated on Voxceleb2.
3.3 Training and Inference
Training. The style-aware lip expert is first pre-trained by determining whether randomly sampled audio and lip motion clips are synchronous as in [49] and then frozen during training the denoising network.

### v2

# DreamTalk: When Emotional Talking Head Generation Meets Diffusion Probabilistic Models

@@ Abstract
Emotional talking head generation has attracted growing attention. Previous methods, which are mainly GAN-based, still struggle to consistently produce satisfactory results across diverse emotions and cannot conveniently specify personalized emotions. In this work, we leverage powerful diffusion models to address the issue and propose DreamTalk, a framework that employs meticulous design to unlock the potential of diffusion models in generating emotional talking heads. Specifically, DreamTalk consists of three crucial components: a denoising network, a style-aware lip expert, and a style predictor. The diffusion-based denoising network can consistently synthesize high-quality audio-driven face motions across diverse emotions. To enhance lip-motion accuracy and emotional fullness, we introduce a style-aware lip expert that can guide lip-sync while preserving emotion intensity. To more conveniently specify personalized emotions, a diffusion-based style predictor is utilized to predict the personalized emotion directly from the audio, eliminating the need for extra emotion reference. By this means, DreamTalk can consistently generate vivid talking faces across diverse emotions and conveniently specify personalized emotions. Extensive experiments validate DreamTalk’s effectiveness and superiority. The code is available at https://github.com/ali-vilab/dreamtalk.
Index Terms:
Emotional talking head generation, Diffusion models
publicationid: pubid:
I Introduction
Audio-driven talking head generation, which concerns animating portraits with speech audio, has garnered significant interest due to its diverse applications, such as film dubbing, digital human generation, video conferences in band-limited conditions, and online education. To produce realistic talking heads, it is crucial to generate and control emotions. Recognizing that a single type of emotion can be expressed in diverse ways, recent research focus has shifted from modeling coarse-grained, discrete emotions to modeling more fine-grained, personalized emotions [1, 2, 3]. These personalized emotions are also called speaking styles [2, 4]. A speaking style is defined as facial motion patterns reflected in a talking video clip. In different video clips, the speakers may have different speaking habits and may display various emotions. Therefore, speaking styles are diverse.
Existing methods still struggle to 1) consistently produce high-quality results across diverse speaking styles and 2) conveniently specify desired speaking styles. Existing methods [3, 5, 2, 6] are mainly based on GANs [7]. GANs’ inherent issues, such as mode collapse and unstable training, impair their performance across diverse speaking styles. Although these methods achieve satisfactory results for a limited range of speaking styles, they struggle with more diverse styles, especially ones unseen during training, often resulting in diminished emotional intensity, inaccurate lip motion, or sudden facial distortions [2]. Another issue is that to specify speaking styles, previous methods often rely on extra references, such as videos [3, 5, 1] or texts [8, 9]. Their acquisition requires extra manual effort and hence is inconvenient.
@@ I Introduction
Figure 1: Leveraging the powerful diffusion models, DreamTalk can consistently generate high-quality talking heads across diverse speaking styles. Furthermore, DreamTalk can conveniently use audio to specify personalized speaking style, obviating the need for additional style references.
@@
As a new line of generative technique, diffusion models [10, 11] have shown capability to produce high-quality results in numerous generative areas [12, 13, 14, 15, 16]. The success of diffusion models, stemming from their superior properties such as powerful distribution learning [12, 17], make them exceptionally promising for exploring emotional talking head generation. However, current diffusion-based talking head approaches [18, 19, 20, 21] primarily concentrate on generating talking heads with neutral expressions or a limited number of discrete emotions, lacking diverse and fine-grained speaking styles. Therefore, exploring the full potential of diffusion models for generating talking heads with diverse speaking styles represents a promising, yet unexplored, research direction.
@@
In this paper, we propose DreamTalk, an emotional talking head generation framework that takes advantage of diffusion models to consistently deliver high performance across diverse speaking styles and reduce the reliance on expensive style references. Specifically, DreamTalk is composed of a denoising network, a style-aware lip expert, and a style predictor. The diffusion-based denoising network produces audio-driven facial motions with the speaking style specified by a reference video. The great distribution-learning property of diffusion models enable the denoising network to consistently produce high-quality results across diverse speaking styles. To enhance the lip-sync, we design a style-aware lip expert that drives the denoising network to produce accurate lip motions under different speaking styles. We observe that previous lip experts, which neglect emotional information, compromise the intensity of generated emotions. To preserve the emotion intensity, we find it important to integrate style information into the lip expert, thereby making it style-aware. Finally, to eliminate the need for additional style references, a diffusion-based style predictor is incorporated to predict personalized speaking styles directly from audio. To predict more personalized emotions, we find it crucial to leverage the correlation between speaker identity and speaking styles; therefore, we provide identity information by incorporating the portrait as input.
The effectiveness of DreamTalk is demonstrated through comprehensive qualitative and quantitative evaluations. DreamTalk can even generate reasonable results for songs in multiple languages, despite these audios being significantly different from those in the training set. In summary, our contributions are as follows:
  • We propose DreamTalk, a diffusion-based framework that can consistently generate talking faces with precise lip-sync as well as rich emotions across diverse speaking styles. We find that diffusion models achieve better results than GANs for more diverse speaking styles.
  • We explore how to use audio alone to predict personalized emotions, making it more convenient than relying on extra videos to specify speaking styles. We discover that incorporating identity information significantly enhances prediction accuracy.
  • We propose a style-aware lip expert that can avoid reducing emotion intensity when providing lip guidance. We find that making the lip expert conditioning on speaking style information is crucial for maintaining emotional fullness.
  • Trained in a classifier-free manner, DreamTalk can use the classifier-free guidance scheme to adjust the intensity of arbitrary speaking styles.
II Related Work
@@
Audio-Driven Talking Head Generation. Audio-driven methods [22, 23, 24, 25, 26, 27] fall into two main categories: person-specific and person-agnostic. Person-specific approaches [28, 29, 30, 31, 32] are constrained to generating videos for speakers seen during training. Many of these [33, 34, 35, 36, 37, 30, 38, 39, 40] first craft 3D facial animations, later converting them into realistic videos. Recent advancements [41, 42, 43, 44, 43, 45] have employed neural radiance fields for modeling, yielding high-fidelity, realistic videos. Conversely, person-agnostic methods [46, 47, 48, 49] target generating videos for unseen speakers. Early methods prioritized lip synchronization [50, 51, 52, 53, 49, 54]. Later works shifted focus to natural facial expressions [19, 55] and head poses [56, 57, 58, 59, 60]. FROND [61] introduces a fine-grained motion model that captures local facial movement keypoints and embeds overall motion context to predict audio-driven facial movements and achieve smooth temporal transitions. However, this method fails to generate emotional expressions during speech, thereby affecting the video’s realism.
@@
Emotional Talking Head Generation. Early methods [9, 6, 62, 63, 64, 31, 30, 65, 66] model expressions in discrete emotions. To model fine-grained emotions, recent methods [5, 3, 2, 67] use an expression reference video and transfer the expressions from that video to the generated one. However, these GAN-based methods cannot consistently achieve high performance across diverse emotions. Our work addresses these issues by using diffusion models.
UniFaceGAN [68] introduces a temporally consistent facial video editing framework that handles both face swapping and face reenactment simultaneously by using a 3D reconstruction model and a novel temporal loss constraint. Facial-Prior-Guided FME Generation [69] enhances facial micro-expression generation by utilizing adaptive weighted prior maps and facial priors to guide motion representation. F3A-GAN [70] employs a 3D geometric flow, termed facial flow, to represent natural facial motion for continuous image synthesis. Although all these methods are related to facial expression generation, none of them can generate accurate lip shapes driven by audio in different emotional contexts.
@@
Conveniently specifying desired speaking styles is also important. Most previous methods rely on reference videos [5, 3, 2] or text [8, 9, 6], which needs human labor. A more user-friendly approach is to derive speaking styles from the input audio. Previous methods can only infer a limited number of discrete emotion classes from audio  [30, 9, 65]. TH-PAD [19] generates expressions only aligned with the audio rhythm, not aligning with the emotional content of the audio. Besides, previous methods neglect information in the input portrait. In this work, we aim to infer personalized emotions using input audio and portraits.
@@
Diffusion Models. Diffusion models [11, 10] have demonstrated strong performance across multiple vision tasks [12, 14, 71, 72, 73, 74], including text-to-image generation [75], human motion generation [17], and video generation [15, 76, 77, 78, 79]. Most diffusion-based methods for talking head generation [18, 80, 81, 82, 83, 84, 19, 85, 86, 87, 88], including EMO [21], AniPortrait [85] and Hallo [89], mainly generate talking heads with neutral emotions and lacks emotional controllability. Besides, the inference of EMO is slow. VASA [20] can only generate a limited number of emotions, lacking diverse, fine-grained speaking styles. In this work, we aim to harness diffusion models for generating and controlling diverse, fine-grained speaking styles in talking heads, presenting a more intricate challenge.
III Method
@@
Figure 2: Illustration of DreamTalk. A style-aware lip expert (b) is first trained to provide lip motion guidance for the denoising network (a). The denoising network is then trained to predict emotional audio-driven face motions. Then, A style predictor (c) is trained to use audio to predict the style code. During inference (d), the speaking style can be specified using style codes that are extracted from videos or derived from audio.
III-A Problem Formulation
@@
Given a portrait $\bm{I}$, a speech $\bm{A}$, and a style reference video $\bm{R}$, our method aims to generate a talking head video with lip motions synchronized with the speech and the speaking style reflected in the reference video. The audio $\bm{A}=[\bm{a}_{i}]_{i=1}^{L}$ is parameterized as a sequence of acoustic features. $\bm{R}$ is a sequence of video frames. The head motions in the generated videos can originate from real videos or be produced by existing methods [60, 4].
@@
Besides, to conveniently specify speaking styles, our method also aims to infer the speaking style using solely the speech and the portrait, obviating the need for extra style references. The inferred speaking style can replace the role of style reference videos in controlling expressions (fig. 1), enabling our method to generate personalized emotions with only speech and portrait.
III-B DreamTalk
@@
As illustrated in fig. 2, DreamTalk comprises 3 key components: a denoising network, a style-aware lip expert, and a style predictor.
@@
The denoising network computes face motion conditioned on the speech and style reference video. The face motion $\bm{M}=[\bm{m}_{l}]_{l=1}^{L}$ is parameterized as a sequence of expression parameters from 3D Morphable Models [90]. The face motion is rendered into video frames by a renderer [91]. The style-aware lip expert provides lip motion guidance under diverse expressions and thus drives the denoising network to achieve accurate lip-sync while preserving emotion fullness. The style predictor can predict the speaking style aligned with that conveyed in speech.
@@
The asterisk(*) indicates quantities that are generated.
@@
The style-aware lip expert encodes the lip motions and audio into respective embeddings conditioned on style reference and then computes the cosine similarity to represent the sync probability. To obtain lip motion information from face motion $\bm{m}$, we first convert $\bm{m}$ into the corresponding face mesh and select vertices in the mouth area as the representation of the lip motion. The lip motion and audio encoders are mainly implemented by MLPs and 1D-convolutions, respectively. The style condition is fused into embeddings by first extracting style features from style reference using a style encoder, which mirrors the architecture of the one in the denoising network but does not share parameters with it, and then concatenating the style features with intermediate feature maps from embedding encoders.
Style Predictor. Specifically, the style predictor $S_{\phi}$ predicts the style code $\bm{s}$ extracted by the style encoder in the trained denoising network. Since we observe that style codes correlate with speaker identity (section IV-D), the style predictor also integrates the portrait as input. The style predictor is instantiated as a diffusion model and is trained to predict the style code itself:
@@
Discussion: Advantages over StyleTalk. Although StyleTalk, a GAN-based baseline, also leverages transformer modules, DreamTalk presents notable advantages: 1) StyleTalk’s modules and loss functions are overly complex, which may cause unstable generation results, while DreamTalk’s are simple, making it more extensible and robust. Since GAN’s mode-collapse issue hampers modeling diverse speaking styles, to enhance emotion intensity, StyleTalk uses a complex dynamic network and up to six loss terms. As discussed in section IV-B, the overly complex design may cause unstable and inferior results. In contrast, DreamTalk, leveraging the power of diffusion models, does not need complex modules and only uses two loss terms, which is much simpler. 2) StyleTalk makes incorrect assumptions about the data, which may impair the performance. To apply losses that enhance emotion intensity, StyleTalk assumes the speaking styles are consistent in a predefined video group. However, as discussed in section IV-D, these speaking styles are actually varied. DreamTalk does not need such an assumption. 3) StyleTalk can only specify speaking styles using videos, while DreamTalk, leveraging the style predictor, can specify styles only using input audio, which is more convenient.
TABLE I: Quantitative Comparisons. We do not receive GC-AVT samples on Voxceleb2. SA is only evaluated on MEAD for emotional methods.
@@
MakeItTalk [55] 0.73/0.57/0.52 0.11/0.24/0.24 3.97/5.12/6.29 5.32/4.55/5.15 2.10/3.16/2.17 -
Wav2Lip [54] 0.80/0.63/0.54 0.18/0.30/0.30 2.72/4.53/5.85 4.05/3.60/4.64 5.26/5.83/5.70 -
PC-AVS [59] 0.50/0.42/0.36 0.07/0.13/0.09 5.83/9.71/12.9 4.97/4.17/7.42 2.18/4.85/4.73 -
AVCT [47] 0.83/0.76/0.64 0.14/0.22/0.23 2.92/2.86/3.62 5.52/3.57/3.71 2.53/4.27/3.89 -
GC-AVT [3] 0.34/0.36/   - 0.14/0.28/   - 8.04/10.2/   - 7.10/6.23/   - 2.42/4.72/   - 18.7
EAMM [5] 0.40/0.40/0.43 0.08/0.14/0.20 6.70/7.03/6.36 6.48/6.86/4.89 1.41/2.54/2.24 20.1
StyleTalk [2] 0.84/0.81/0.66 0.16/0.30/0.29 2.12/1.96/2.92 3.25/2.41/2.96 3.47/4.82/4.51 74.2
SadTalker [60] 0.69/0.77/0.44 0.16/0.24/0.19 4.12/5.99/9.12 4.37/4.07/6.11 2.76/4.35/4.38 -
PD-FGC [1] 0.49/0.41/0.35 0.05/0.13/0.12 5.50/9.50/12.5 4.10/4.23/8.19 2.27/4.68/4.64 43.9
EAT [6] 0.53/0.59/0.47 0.15/0.26/0.20 5.54/3.86/5.53 4.79/4.03/5.88 2.16/4.54/4.35 36.4
DreamTalk 0.86/0.85/0.69 0.16/0.31/0.30 1.93/1.80/2.69 2.91/2.15/2.72 3.78/5.17/4.90 86.7
Ground Truth 1/1/1 0.22/0.31/0.33 0/0/0 0/0/0 4.13/5.44/5.23 92.5
III-C Training and Inference
Training. The style-aware lip expert is first pre-trained by determining whether randomly sampled audio and lip motion clips are synchronous as in [54] and then frozen during training the denoising network. We use cosine-similarity with binary cross-entropy loss during training. Specifically, we compute cosine-similarity for the face motion embedding $\bm{e}^{m}$ and audio embedding $\bm{e}^{a}$ to represent the probability that the input audio-motion pair is synchronized. The training loss of the lip expert is:
$\mathcal{L}_{{\text{expert}}}={\text{BCE}}(\frac{\bm{e}^{m}\cdot\bm{e}^{a}}{\max(||\bm{e}^{m}||_{2}\cdot||\bm{e}^{a}||_{2})}).$ (5)